# Xinxu Wei1       Kanhao Zhao1       Yong Jiao1       Lifang He1       Yu Zhang [^1] footnotemark: 1

### v1

@@ 5 Conclusion
We propose BrainGFM, a graph-based brain foundation model pre-trained on heterogeneous fMRI brain graphs constructed from diverse atlases and parcellation schemes. To enhance its generalization and adaptability, we introduce a meta-learning framework to optimize graph prompts, enabling robust few-shot learning under limited data. In addition, we incorporate language prompt tokens to guide zero-shot generalization, allowing BrainGFM to transfer effectively across unseen datasets, tasks, atlases, and neurological disorders. Our large-scale, multi-atlas fMRI dataset provides a rich and diverse training corpus, and BrainGFM demonstrates superior performance in both effectiveness and efficiency compared to prior time-series-based and ROI-based foundation models.
@@ Appendix B Contributions of BrainGFM for Unifying Cohorts, Atlases, and Disorders
Appendix C Comparison of Our Vanilla Graph FM and BrainGFM with Prior Time-Series-based and ROI-based Brain FMs
@@
Appendix D Ablation Study on Different Tuning Methods
As shown in Table 3, after completing model pre-training, we explored three downstream adaptation strategies: full fine-tuning Sun et al. [2022], parameter-efficient fine-tuning (PEFT) Ding et al. [2023], and graph prompt-tuning Sun et al. [2023]. Full fine-tuning updates all model parameters during downstream training, offering strong performance but at a high computational cost. In contrast, PEFT methods reduce training overhead by modifying only a small subset of parameters or introducing lightweight modules. We specifically evaluate two popular PEFT variants: prefix-tuning and LoRA. Note that the details of all tuning methods are summarized in Table 4.
@@ Appendix D Ablation Study on Different Tuning Methods
Appendix E Brain fMRI Graph Construction from fMRI Time Series
@@
Appendix F Comparison of Positional Encoding Methods
We investigate three commonly used positional encoding (PE) strategies for graph neural networks: Laplacian Positional Encoding (LPE), Node Degree Positional Encoding, and Random Walk Structural Encoding (RWSE). Below, we define each method and evaluate their characteristics in the context of fMRI-based brain graph modeling.
@@ Appendix F Comparison of Positional Encoding Methods
[PE] Type Pre-Trained Efficiency Performance (ACC / AUC) w/o [PE] Very Fast 65.2 / 67.1 w/o [PE] Very Fast 69.5 / 71.7 Laplacian [PE] Slow 69.2 / 71.3 Node Degree [PE] Fast 68.4 / 70.5 RWSE [PE] Fast 66.1 / 68.0 RWSE [PE] Fast 70.5 / 73.3
Appendix G Unified Training on Brain Graphs with Varying Numbers of Nodes
@@
Appendix H Graph Prompt Construction and Insertion
@@
Appendix I Meta-Learning for Unifying Diverse Atlases and Disorders
@@
Appendix J Details of Graph Masked Autoencoder (GMAE) and Graph Contrastive Learning (GCL)
@@
J.1 Graph Masked Autoencoder (GMAE)
Inspired by recent advances in masked autoencoding Hou et al. [2022], we apply random masking to both nodes and edges of the input graph. Given a graph with $N$ nodes, adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, and node features $\mathbf{X}\in\mathbb{R}^{N\times F}$, we randomly sample a subset of nodes $\mathcal{V}_{M}\subset\mathcal{V}$ to mask. For each masked node $v_{i}\in\mathcal{V}_{M}$, its input feature is replaced with a learnable mask token $\mathbf{x}_{[M]}\in\mathbb{R}^{D}$. The masked node feature matrix $\tilde{\mathbf{X}}$ is defined as:
@@
where $\mathbf{M}_{e}\in\{0,1\}^{N\times N}$ is a symmetric binary mask sampled from a Bernoulli distribution with parameter $1-p$, and $\odot$ denotes element-wise multiplication.
@@
J.2 Graph Contrastive Learning (GCL)
In parallel with the generative pathway, we apply contrastive learning to enforce view-invariant representations. Specifically, we generate two augmented views of the same input graph by applying lightweight stochastic perturbations (i.e., random feature dropout and edge dropout). One view is treated as the query, while the other serves as the key. Let $(\mathbf{X}^{(q)},\mathbf{A}^{(q)})$ and $(\mathbf{X}^{(k)},\mathbf{A}^{(k)})$ denote the two views; these are passed through a shared encoder to obtain corresponding representations $z_{q}$ and $z_{k}$. We adopt the NT-Xent contrastive loss Qiu et al. [2020] to maximize the similarity between matching query-key pairs from the same graph while distinguishing them from others in the batch:
@@
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, $\tau$ is a temperature hyperparameter, and $B$ is the batch size. Here, $b$ and $b^{\prime}$ index different samples within the batch, where each sample corresponds to an augmented graph.
@@
Appendix K Comparison between Two Graph Pre-Training Methods
@@
Appendix L Implement Details
During pre-training, we set the batch size to 128 and used the Adam optimizer with a learning rate of 0.0001. The number of training epochs was set to 100. For downstream classification tasks, we set the batch size to 16 in the full-shot setting, and to 1 in both the few-shot and zero-shot settings. The Adam optimizer was also used for these tasks, with the same learning rate of 0.0002. In the meta-learning setup, we trained the model for 30 epochs. More settings about pre-training and graph transformer backbone and meta learning can be found in Table 7, 8, 10, 9.
@@
Appendix M Baselines
The Table 11 provides a systematic comparison of various brain foundation models and baseline methods across multiple dimensions, including architectural type, data domain, pre-training strategy, and tuning method. These approaches can be broadly categorized into three groups: conventional models without pre-training (e.g., Vanilla GCN Kipf and Welling [2016], BrainGNN Li et al. [2021], Vanilla ROI-based Tansformer (TF) Yun et al. [2019]), ROI- or time-series-based pre-trained models (e.g., BrainNPT Hu et al. [2024], BrainMass Yang et al. [2024], BrainLM Caro et al. [2023]), and our proposed graph-based foundation model, BrainGFM.
The conventional models do not leverage any pre-training and are limited to single parcellation and single-disorder settings, resulting in restricted generalization capabilities. ROI-based models such as BrainNPT and BrainMass employ generative pre-training on region-level features, enabling improved performance across a limited number of disorders. BrainLM further enhances temporal modeling through time-series-based generative pre-training.
@@
Model Foundation Model Type Domain Pre-Training Method Tuning Method Tuning Shot Parcellation Disorder Vanilla GCN Graph Spatial - - - Single Single BrainGNN Graph Spatial - - - Single Single Vanilla TF ROI Spatial - - - Single Single Graph TF Graph Spatial - - - Single Single BrainNetTF ROI Spatial - - - Single Single BrainNPT ROI Spatial ROI Generative Fine-Tuning Full/Few-shot Single Multiple (< 5) BrainMass ROI Spatial ROI Generative Fine-Tuning Full/Few-shot Single Multiple (10) BrainLM Time Series Temporal Time Series Generative Fine-Tuning Full-shot Single Multiple (< 5) BrainGFM Graph Spatial Graph (Gener. + Contra.) Fine/Prompt-Tuning Full/Few/Zero-shot Multiple Multiple (25)
Appendix N Benchmarks, Datasets, Disorders and Downstream Tasks
@@
Pre-Train ABCD Casey et al. [2018] 11,878 35,770 Multiple
ADHD 200 consortium [2012] 973 1382 ADHD
ABIDE I Di Martino et al. [2014] 1,112 1,112 ASD
ADNI 3 Jack Jr et al. [2008] 1,071 1,410 AD, MCI
AOMIC Snoek et al. [2021] 210 210 Multiple
AURORA McLean et al. [2020] 284 284 PTSD
CAM_CAN Shafto et al. [2014] 652 652 -
CATD Nielson et al. [2023] 127 454 Multiple
GSP Holmes et al. [2015] 1,569 2,706 -
HCP-Aging Bookheimer et al. [2019] 724 724 -
EMBARC Trivedi et al. [2016] 308 308 MDD
LEMON Babayan et al. [2019] 213 213 MDD
HABS Dagley et al. [2017] 284 1,371 -
PREVEND_AD Tremblay-Mercier et al. [2021] 343 2,427 AD
SRPBS_Japan Yamashita et al. [2019] 1,410 1,410 ASD
NYU_CUD Kelly et al. [2011] 29 56 CUD
OASIS3 LaMontagne et al. [2019] 1,172 4,090 Dementia
HBN Alexander et al. [2017] 2,228 4,039 Multiple
SubMex_RTMS Angeles-Valdez et al. [2024] 150 150 CUD
Internal Test ADHD 200 consortium [2012] 973 1382 ADHD
HBN Alexander et al. [2017] 2,282 4,039 Multiple
OASIS3 LaMontagne et al. [2019] 1,172 4,090 Dementia
Semi-External Test ABIDE II Di Martino et al. [2014] 1,044 1,044 ASD
ADNI 2 Weiner et al. [2013] 1,171 1,306 AD, MCI
SubMex_CUD Angeles-Valdez et al. [2022] 135 135 CUD
External Test UCLA_CNP Poldrack et al. [2016] 261 261 Multiple
REST-META-MDD Yan et al. [2019] 2,379 2,379 MDD
Table 13 and 14 summarizes the 25 brain disorder classification tasks selected from various public neuroimaging datasets. These tasks span a remarkably broad range in terms of disease types, age groups, and data sources, reflecting the diversity and complexity of real-world clinical scenarios. The downstream evaluation covers a wide spectrum of brain conditions, including neurodevelopmental disorders (e.g., ADHD, ASD, SLD), affective and emotional disorders (e.g., MDD, anxiety, bipolar disorder), psychotic disorders (e.g., schizophrenia), neurodegenerative diseases (e.g., AD, MCI, dementia), and substance use disorders (e.g., CUD). Subject age ranges from as young as 5 years old (e.g., ABIDE II, HBN) to elderly adults nearing 90 years old (e.g., ADNI, OASIS3), capturing the full human lifespan from brain development to cognitive decline. To ensure scientific rigor and fairness, we carefully constructed sex-balanced subsets for all labeled downstream tasks, meaning that each task includes approximately equal numbers of male and female samples. This prevents potential sex biases from influencing model performance. For multi-diagnostic datasets like HBN, we created multiple binary classification tasks (e.g., MDD vs. NC, ADHD vs. NC), each treated independently within a unified evaluation framework. By incorporating such a comprehensive and diverse benchmark, we are able to thoroughly assess the robustness, transferability, and clinical relevance of our proposed BrainGFM.
@@
Dataset Disease/Disorder/Disability Downstream Task Age Sample Size ADHD200 Att-ention-Deficit/Hyperactivity Disorder (ADHD) ADHD vs. NC 8-26 402/580 ABIDE II Autistic Spectrum Disorders (ASD) ASD vs. NC 5-64 581/733 ADNI 2 Alzheimers Disease (AD) AD vs. NC 55-89 91/100 Mild Cognitive Impairment (MCI) MCI vs. NC 55-89 168/200 OASIS3 Dementia (DM) DM vs. NC 45-88 290/300 HBN Major Depression Disorder (MDD) MDD vs. NC 6-20 261/310 Anxiety (ANX) ANX vs. NC 6-20 224/250 Oppositional Defiant Disorder (ODD) ODD vs. NC 6-20 519/319 Obsessive-Compulsive Disorder (OCD) OCD vs. NC 6-20 130/150 Language Disorder (LD) LD vs. NC 6-20 464/319 Specific Learning Disorder (SLD) SLD vs. NC 6-20 335/319 Enuresis (ENU) ENU vs. NC 6-20 279/319 Intellectual Disability (ID) ID vs. NC 6-20 129/150 Post-Traumatic Stress Disorder (PTSD) PTSD vs. NC 6-20 39/40 Encopresis (ECP) ECP vs. NC 6-20 68/70 Dysthymia (PDD) PDD vs. NC 6-20 85/90 Tourette Sydnrome (PS) PS vs. NC 6-20 68/70 Adjustment Disorder (AJD) AJD vs. NC 6-20 96/100 Provisional Tic Disorder (PTD) PTD vs. NC 6-20 68/70 Motor Disorder (MD) MD vs. NC 6-20 123/130 Speech Sound Disorder (SSD) SSD vs. NC 6-20 98/100 Communication Disorder (CD) CD vs. NC 6-20 43/45 SubMex_CUD Cocaine Use Disorder (CUD) CUD vs. NC 18-45 72/63 UCLA_CNP Schizophrenia (SCHZ) SCHZ vs. NC 21-50 50/55 Bipolar Disorder (BP) BP vs. NC 21-50 49/55 REST-META-MDD Major Depression Disorder (MDD) MDD vs. NC 21-50 1,252/1,101
Appendix O Data Acquisition and Preprocessing
@@
For all cohorts, resting-state fMRI data were collected with varying protocols and scanner parameters specific to each study site. All available resting-state fMRI data were preprocessed using the well-established fMRIPrep pipeline Esteban et al. [2019]. The T1-weighted image was corrected for intensity non-uniformity and then stripped skull. Spatial normalization was done through nonlinear registration, with the T1w reference Avants et al. [2008]. Using FSL, brain features such as cerebrospinal fluid, white matter, and grey matter were segmented from the reference, brain-extracted T1 weighted image Zhang et al. [2000]. The fieldmap information was used to correct distortion in low-frequency and high-frequency components of fieldmap. Then, a corrected echo-planar imaging reference was obtained from a more accurate co-registration with the anatomical reference. The blood-oxygenation-level-dependent (BOLD) reference was then transformed to the T1-weighted image with a boundary-based registration method, configured with nine degrees of freedom to account for distortion remaining in the BOLD reference Greve and Fischl [2009]. Head-motion parameters (rotation and translation parameters of volume-to-reference transform matrices) were estimated with MCFLIRT (FSL). BOLD signals were slice-time corrected and resampled onto the participant’s original space with head-motion correction, susceptibility distortion’s correction, and then resampled into standard space, generating a preprocessed BOLD run in MNI152NLin2009cAsym space. Automatic removal of motion artifacts using independent component analysis (ICA-AROMA) Pruim et al. [2015] was performed on the preprocessed BOLD time-series on MNI space after removal of non-steady-state volumes and spatial smoothing with an isotropic Gaussian kernel of 6 mm FWHM (full-width half-maximum).
@@
Appendix P Definition and Description of Used Atlases/Parcellations
@@
Appendix Q Broader Impact

### v2

@@ 5 Conclusion
We propose BrainGFM, a graph-based brain foundation model pre-trained on heterogeneous fMRI brain graphs constructed from diverse atlases and parcellation schemes. To enhance its generalization and adaptability, we introduce a meta-learning framework to optimize graph prompts, enabling robust few-shot learning under limited data. In addition, we incorporate language prompt tokens to guide zero-shot generalization, allowing BrainGFM to transfer effectively across unseen datasets, tasks, atlases, and neurological disorders. Our large-scale, multi-atlas fMRI dataset provides a rich and diverse training corpus, and BrainGFM demonstrates superior performance in both effectiveness and efficiency compared to prior time-series-based and Connectome/FC-based foundation models.
@@ Appendix B Contributions of BrainGFM for Unifying Cohorts, Atlases, and Disorders
Appendix C Comparison of Our Vanilla Graph FM and BrainGFM with Prior Time-Series-based and Connectome/FC-based Brain FMs
@@
Appendix D Ablation Study for Full/Few/Zero-Shot on Graph/Language Prompt and Meta Learning on Unseen Disorders
Figure 9: Performance comparison on two uncommon disorders from the HBN dataset under Full-Shot, Few-Shot, and Zero-Shot settings. The HBN dataset is excluded from the pre-training stage to ensure that the tested disorders are unseen during training. The results highlight the effectiveness of incorporating graph prompts (G-Prompt), meta-learning (Meta L.), and language prompts (Lan. Prompt) into the BrainGFM model, particularly in few-shot and zero-shot scenarios.
Appendix E Ablation Study for Different Brain Graph Construction Methods
[…]
Corpus Top-K Correlation KNN ABIDE II ASD 73.5 73.9 HBN MDD 85.5 86.7
Appendix F Cross-Atlas/Parcellation Experiments
Table 4: Ablation experiments on the ADNI2 dataset for Alzheimer’s disease (AD) classification. The reported improvements (%) indicate the performance gain over models without pre-training.
Type Pre-training Fine-tuning Performance
Cross-Atlas Schaefer100 AAL116 3.3% ±1.5% $\uparrow$
Schaefer100 Power264 2.7% ±1.3% $\uparrow$
Schaefer200 + AAAL116 Schaefer100 3.2% ±1.6% $\uparrow$
Cross-Parcellation Schaefer100 Schaefer300 2.7% ±1.4% $\uparrow$
Schaefer100 + Schaefer300 Schaefer200 3.5% ±1.5% $\uparrow$
AAAL116 AAL3v1 3.1% ±1.3% $\uparrow$
Appendix G Ablation Study on Different Tuning Methods
As shown in Table 5, after completing model pre-training, we explored three downstream adaptation strategies: full fine-tuning [Sun et al., 2022], parameter-efficient fine-tuning (PEFT) [Ding et al., 2023], and graph prompt-tuning [Sun et al., 2023]. Full fine-tuning updates all model parameters during downstream training, offering strong performance but at a high computational cost. In contrast, PEFT methods reduce training overhead by modifying only a small subset of parameters or introducing lightweight modules. We specifically evaluate two popular PEFT variants: prefix-tuning and LoRA. Note that the details of all tuning methods are summarized in Table 6.
@@ Appendix G Ablation Study on Different Tuning Methods
Appendix H Brain fMRI Graph Construction from fMRI Time Series
@@
Appendix I Comparison of Positional Encoding Methods
We investigate four commonly used positional encoding (PE) strategies for graph neural networks: Laplacian Positional Encoding (LPE), Node Degree Positional Encoding, Brain Gradient Positional Encoding [Dong et al., 2024a], and Random Walk Structural Encoding (RWSE). Below, we define each method and evaluate their characteristics in the context of fMRI-based brain graph modeling.
@@ Appendix I Comparison of Positional Encoding Methods
[PE] Type Pre-Trained Efficiency Performance (ACC / AUC) w/o [PE] Very Fast 65.2 / 67.1 w/o [PE] Very Fast 69.5 / 71.7 Laplacian [PE] Slow 69.2 / 71.3 Node Degree [PE] Fast 68.4 / 70.5 Brain Gradient [PE] Fast 70.4 / 72.5 RWSE [PE] Fast 66.1 / 68.0 RWSE [PE] Fast 70.5 / 73.3
Appendix J Unified Training on Brain Graphs with Varying Numbers of Nodes
@@
Appendix K Graph Prompt Construction and Insertion
@@
Appendix L Meta-Learning for Unifying Diverse Atlases and Disorders
@@
Appendix M Comparison of our Meta-learning with the Meta-Matching
Our method adopts a prompt-based meta-learning strategy to enable efficient transfer across neuroimaging tasks defined by different disorders and parcellation schemes. Specifically, we introduce learnable prompt tokens that serve as task-specific parameters injected into the backbone model. These prompts are meta-trained through episodic learning with inner- and outer-loop optimization, allowing the model to adapt to new tasks without updating the backbone. This design enables parameter-efficient adaptation and supports both few-shot and zero-shot generalization, as the prompts can be tuned or inferred for unseen tasks based on their task identity.
A related concept appears in Meta-Matching [He et al., 2022], which aims to transfer pre-trained models from large-scale datasets (e.g., UK Biobank) to small-scale neuroimaging datasets. Meta-Matching freezes the entire backbone and learns a lightweight linear regressor on the model outputs to align with new target labels. Each target dataset is treated as an independent task, without leveraging shared structure or task distributions. While both approaches seek efficient cross-dataset transfer, their adaptation mechanisms and learning paradigms are fundamentally different.
As summarized in Tables 8 and 9, both methods address domain shift and operate in low-resource settings by introducing parameter-efficient task modules. However, unlike Meta-Matching’s post-hoc linear adaptation, our method performs task-level meta-learning by explicitly optimizing prompt tokens via backpropagation. This allows our model to generalize not only to few-shot settings but also to zero-shot tasks, making it more flexible and extensible across diverse brain graph domains.
Table 8: Conceptual Similarities between Meta-Matching and Our Prompt-based Meta-Learning Framework
Aspect Shared Property Goal Both aim to transfer predictive models from large-scale to small-scale neuroimaging datasets. Data Efficiency Both methods improve learning in low-resource settings without retraining the entire backbone. Parameter Efficiency Both adopt parameter-efficient adaptation by freezing the main model and training small modules. Domain Shift Adaptation Both address distributional shifts across cohorts, sites, or acquisition protocols. Task-specific Specialization Both customize predictions for downstream tasks using lightweight task-specific components.
Table 9: Key Differences between Meta-Matching [He et al., 2022] and Our Prompt-based Meta-Learning Framework
Aspect Meta-Matching Our Meta-Learning Framework Meta-learning Paradigm No explicit meta-learning; post-hoc linear adaptation Prompt-based meta-learning with inner/outer optimization Adaptation Mechanism Learns output mapping $g(f_{\text{pre}}(x))$ via linear regressor Learns task-specific prompt tokens injected into the backbone model Learnable Modules Linear output head trained per task Prompt tokens trained per task and meta-updated across tasks Gradient Flow Only regressor is trained; backbone remains fixed Prompts are updated via inner-loop gradients; backbone frozen Task Definition Each small dataset = separate task; no shared structure Tasks defined by (disorder, atlas) tuples sampled from unified distribution Update Scope No model component is updated during adaptation Only prompt modules are adapted; base encoder remains fixed Generalization Scope Few-shot only; does not support zero-shot transfer Supports both few-shot and zero-shot generalization via prompt inference
Appendix N Details of Graph Masked Autoencoder (GMAE) and Graph Contrastive Learning (GCL)
@@
N.1 Graph Masked Autoencoder (GMAE)
Inspired by recent advances in masked autoencoding [Hou et al., 2022], we apply random masking to both nodes and edges of the input graph. Given a graph with $N$ nodes, adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, and node features $\mathbf{X}\in\mathbb{R}^{N\times F}$, we randomly sample a subset of nodes $\mathcal{V}_{M}\subset\mathcal{V}$ to mask. For each masked node $v_{i}\in\mathcal{V}_{M}$, its input feature is replaced with a learnable mask token $\mathbf{x}_{[M]}\in\mathbb{R}^{D}$. The masked node feature matrix $\tilde{\mathbf{X}}$ is defined as:
@@
[…]
@@
N.2 Graph Contrastive Learning (GCL)
In parallel with the generative pathway, we apply contrastive learning to enforce view-invariant representations. Specifically, we generate two augmented views of the same input graph by applying lightweight stochastic perturbations (i.e., random feature dropout and edge dropout). One view is treated as the query, while the other serves as the key. Let $(\mathbf{X}^{(q)},\mathbf{A}^{(q)})$ and $(\mathbf{X}^{(k)},\mathbf{A}^{(k)})$ denote the two views; these are passed through a shared encoder to obtain corresponding representations $z_{q}$ and $z_{k}$. We adopt the NT-Xent contrastive loss [Qiu et al., 2020] to maximize the similarity between matching query-key pairs from the same graph while distinguishing them from others in the batch:
@@
[…]
@@
Appendix O Comparison between Two Graph Pre-Training Methods
@@
Appendix P Implement Details
During pre-training, we set the batch size to 128 and used the Adam optimizer with a learning rate of 0.0001. For downstream classification tasks, we set the batch size to 16 in the full-shot setting, and to 1 in both the few-shot and zero-shot settings. The Adam optimizer was also used for these tasks, with the same learning rate of 0.0002. In the meta-learning setup, we trained the model for 50 epochs. More settings about pre-training and graph transformer backbone and meta learning can be found in Table 11, 12, 14, 13.
@@
Appendix Q Baselines
Model Foundation Input Type Domain Pre-Training Method Tuning Method Tuning Shot Parcellation Disorder Vanilla GCN Graph Spatial - - - Single Single BrainGNN Graph Spatial - - - Single Single Vanilla TF Connectome/FC Spatial - - - Single Single Graph TF Graph Spatial - - - Single Single BrainNetTF Connectome/FC Spatial - - - Single Single BrainNPT Connectome/FC Spatial Connectome/FC Generative Fine-Tuning Full/Few-shot Single Multiple (< 5) BrainMass Connectome/FC Spatial Connectome/FC Generative Fine-Tuning Full/Few-shot Single Multiple (10) BrainLM Time Series Temporal Time Series Generative Fine-Tuning Full-shot Single Multiple (< 5) Brain-JEPA Time Series Spatial+Temporal Time Series Generative Fine-Tuning Full-shot Single Multiple (< 5) BrainGFM Graph Spatial Graph (Gener. + Contra.) Fine/Prompt-Tuning Full/Few/Zero-shot Multiple Multiple (25)
The Table 15 provides a systematic comparison of various brain foundation models and baseline methods across multiple dimensions, including architectural type, data domain, pre-training strategy, and tuning method. These approaches can be broadly categorized into three groups: conventional models without pre-training (e.g., Vanilla GCN [Kipf and Welling, 2016], BrainGNN [Li et al., 2021], Vanilla Connectome/FC-based Tansformer (TF) [Yun et al., 2019]), ROI- or time-series-based pre-trained models (e.g., BrainNPT [Hu et al., 2024], BrainMass [Yang et al., 2024], BrainLM [Caro et al., 2023]), and our proposed graph-based foundation model, BrainGFM.
The conventional models do not leverage any pre-training and are limited to single parcellation and single-disorder settings, resulting in restricted generalization capabilities. Connectome/FC-based models such as BrainNPT and BrainMass employ generative pre-training on region-level features, enabling improved performance across a limited number of disorders. BrainLM further enhances temporal modeling through time-series-based generative pre-training.
@@
Appendix R Benchmarks, Datasets, Disorders and Downstream Tasks
@@
Pre-Train ABCD [Casey et al., 2018] 11,878 35,770 Multiple
ADHD 200 [consortium, 2012] 973 1382 ADHD
ABIDE I [Di Martino et al., 2014] 1,112 1,112 ASD
ADNI 3 [Jack Jr et al., 2008] 1,071 1,410 AD, MCI
AOMIC [Snoek et al., 2021] 210 210 Multiple
AURORA [McLean et al., 2020] 284 284 PTSD
CAM_CAN [Shafto et al., 2014] 652 652 -
CATD [Nielson et al., 2023] 127 454 Multiple
GSP [Holmes et al., 2015] 1,569 2,706 -
HCP-Aging [Bookheimer et al., 2019] 724 724 -
EMBARC [Trivedi et al., 2016] 308 308 MDD
LEMON [Babayan et al., 2019] 213 213 MDD
HABS [Dagley et al., 2017] 284 1,371 -
PREVEND_AD [Tremblay-Mercier et al., 2021] 343 2,427 AD
SRPBS_Japan [Yamashita et al., 2019] 1,410 1,410 ASD
NYU_CUD [Kelly et al., 2011] 29 56 CUD
OASIS3 [LaMontagne et al., 2019] 1,172 4,090 Dementia
HBN [Alexander et al., 2017] 2,228 4,039 Multiple
SubMex_RTMS [Angeles-Valdez et al., 2024] 150 150 CUD
Internal Test ADHD 200 [consortium, 2012] 973 1382 ADHD
HBN [Alexander et al., 2017] 2,282 4,039 Multiple
OASIS3 [LaMontagne et al., 2019] 1,172 4,090 Dementia
Semi-External Test ABIDE II [Di Martino et al., 2014] 1,044 1,044 ASD
ADNI 2 [Weiner et al., 2013] 1,171 1,306 AD, MCI
SubMex_CUD [Angeles-Valdez et al., 2022] 135 135 CUD
External Test UCLA_CNP [Poldrack et al., 2016] 261 261 Multiple
REST-META-MDD [Yan et al., 2019] 2,379 2,379 MDD
Table 17 and 18 summarizes the 25 brain disorder classification tasks selected from various public neuroimaging datasets. These tasks span a remarkably broad range in terms of disease types, age groups, and data sources, reflecting the diversity and complexity of real-world clinical scenarios. The downstream evaluation covers a wide spectrum of brain conditions, including neurodevelopmental disorders (e.g., ADHD, ASD, SLD), affective and emotional disorders (e.g., MDD, anxiety, bipolar disorder), psychotic disorders (e.g., schizophrenia), neurodegenerative diseases (e.g., AD, MCI, dementia), and substance use disorders (e.g., CUD). Subject age ranges from as young as 5 years old (e.g., ABIDE II, HBN) to elderly adults nearing 90 years old (e.g., ADNI, OASIS3), capturing the full human lifespan from brain development to cognitive decline. To ensure scientific rigor and fairness, we carefully constructed sex-balanced subsets for all labeled downstream tasks, meaning that each task includes approximately equal numbers of male and female samples. This prevents potential sex biases from influencing model performance. For multi-diagnostic datasets like HBN, we created multiple binary classification tasks (e.g., MDD vs. NC, ADHD vs. NC), each treated independently within a unified evaluation framework. By incorporating such a comprehensive and diverse benchmark, we are able to thoroughly assess the robustness, transferability, and clinical relevance of our proposed BrainGFM.
@@
Dataset Disease/Disorder/Disability Downstream Task Age Sample Size ADHD200 Att-ention-Deficit/Hyperactivity Disorder (ADHD) ADHD vs. NC 8-26 402/580 ABIDE II Autistic Spectrum Disorders (ASD) ASD vs. NC 5-64 581/733 ADNI 2 Alzheimers Disease (AD) AD vs. NC 55-89 91/100 Mild Cognitive Impairment (MCI) MCI vs. NC 55-89 168/200 OASIS3 Dementia (DM) DM vs. NC 45-88 290/300 HBN Major Depression Disorder (MDD) MDD vs. NC 6-20 261/100 Anxiety (ANX) ANX vs. NC 6-20 224/250 Oppositional Defiant Disorder (ODD) ODD vs. NC 6-20 519/319 Obsessive-Compulsive Disorder (OCD) OCD vs. NC 6-20 130/150 Language Disorder (LD) LD vs. NC 6-20 464/319 Specific Learning Disorder (SLD) SLD vs. NC 6-20 335/319 Enuresis (ENU) ENU vs. NC 6-20 279/319 Intellectual Disability (ID) ID vs. NC 6-20 129/150 Post-Traumatic Stress Disorder (PTSD) PTSD vs. NC 6-20 39/40 Encopresis (ECP) ECP vs. NC 6-20 68/70 Dysthymia (PDD) PDD vs. NC 6-20 85/90 Tourette Sydnrome (TS) PS vs. NC 6-20 68/70 Adjustment Disorder (AJD) AJD vs. NC 6-20 96/100 Provisional Tic Disorder (PTD) PTD vs. NC 6-20 68/70 Motor Disorder (MD) MD vs. NC 6-20 123/130 Speech Sound Disorder (SSD) SSD vs. NC 6-20 98/100 Communication Disorder (CD) CD vs. NC 6-20 43/45 SubMex_CUD Cocaine Use Disorder (CUD) CUD vs. NC 18-45 72/63 UCLA_CNP Schizophrenia (SCHZ) SCHZ vs. NC 21-50 50/55 Bipolar Disorder (BP) BP vs. NC 21-50 49/55 REST-META-MDD Major Depression Disorder (MDD) MDD vs. NC 21-50 1,252/1,101
Appendix S Data Acquisition and Preprocessing
@@
For all cohorts, resting-state fMRI data were collected with varying protocols and scanner parameters specific to each study site. All available resting-state fMRI data were preprocessed using the well-established fMRIPrep pipeline [Esteban et al., 2019]. The T1-weighted image was corrected for intensity non-uniformity and then stripped skull. Spatial normalization was done through nonlinear registration, with the T1w reference [Avants et al., 2008]. Using FSL, brain features such as cerebrospinal fluid, white matter, and grey matter were segmented from the reference, brain-extracted T1 weighted image [Zhang et al., 2000]. The fieldmap information was used to correct distortion in low-frequency and high-frequency components of fieldmap. Then, a corrected echo-planar imaging reference was obtained from a more accurate co-registration with the anatomical reference. The blood-oxygenation-level-dependent (BOLD) reference was then transformed to the T1-weighted image with a boundary-based registration method, configured with nine degrees of freedom to account for distortion remaining in the BOLD reference [Greve and Fischl, 2009]. Head-motion parameters (rotation and translation parameters of volume-to-reference transform matrices) were estimated with MCFLIRT (FSL). BOLD signals were slice-time corrected and resampled onto the participant’s original space with head-motion correction, susceptibility distortion’s correction, and then resampled into standard space, generating a preprocessed BOLD run in MNI152NLin2009cAsym space. Automatic removal of motion artifacts using independent component analysis (ICA-AROMA) [Pruim et al., 2015] was performed on the preprocessed BOLD time-series on MNI space after removal of non-steady-state volumes and spatial smoothing with an isotropic Gaussian kernel of 6 mm FWHM (full-width half-maximum).
@@
Appendix T Definition and Description of Used Atlases/Parcellations
@@
Appendix U Broader Impact